OpenStudy (anonymous):

Transversal waves pass with a speed of 150 m/s through an 80 cm wire with 550 N tension. Which is the wires mass?

OpenStudy (shane_b):

\[v=\sqrt{\frac{TL}{M}}\]\[150m/s=\sqrt{\frac{(550N)(0.80m)}{M}}\]Now just solve for M.